# New Hardness Results for the LOCAL Model via a Simple Self-Reduction

### v1

@@ 1.1 Lower bound techniques in the LOCAL model
On a high level, lower bounds shown via the round elimination technique consist in constructing a sequence of problems that satisfy some desired properties. More in detail, let $P_{0}$ be a problem of interest for which we want to show a lower bound. The goal (and the challenge) is to produce a sequence $P_{0},P_{1},\ldots,P_{T}$ of problems such that the following two properties hold: (i) for all $0<i\leq T$ it holds that $P_{i}$ is one round easier that $P_{i-1}$; (ii) $P_{T}$ cannot be solved in 0 rounds. Such a sequence implies that $P_{0}$ cannot be solved in $T$ rounds and hence we obtain a lower bound for $P_{0}$. The first LOCAL lower bounds based on this technique date back to the late ’80s and the beginning of the ’90s, and they show that $3$-coloring cycles requires $\Omega(\operatorname{log}^{*}n)$ rounds (where $n$ is the number of nodes in the graph), both for deterministic [Lin92] and randomized [Nao91] algorithms. Even though at a first glance such a technique may look conceptually easy, for many years it was very challenging for researchers to use it for showing lower bounds for different graph problems in the LOCAL model. In fact, the first round elimination lower bound after the aforementioned ones was shown more than 25 years later, in a paper that presents a lower bound for the sinkless orientation problem, [^1] A solution to sinkless orientation requires orienting edges such that no node is a sink. the proof of which is very long and involved [BFH+16].
@@
Lower bounds shown via indistinguishability arguments consist in constructing a class of graphs that have some desired properties. On a very high level, this technique consists in showing that there are nodes that, within their running time, cannot distinguish between two different settings in which they would be required to give different outputs, reaching a contradiction. Perhaps one of the most influential indistinguishability-type lower bounds is the so-called KMW lower bound, shown in 2004, which provides a lower bound for many fundamental graph problems [KMW04]. The KMW construction is very nice and clever. However, the proof is quite complex. For more than 15 years, the KMW construction was one of the few (non-trivial) indistinguishability-type lower bounds, and in fact, until recently, there were no new constructions inspired by the KMW one. Arguably, the involved proof of the KMW lower bound may have been an obstacle.
@@ 1.2 Simplifying proofs and techniques marked a turning point
Regarding lower bounds proved via the round elimination technique, the turning point happened in 2019, when the automatic round elimination technique was introduced [Bra19]. Before this work, we had to guess what could be the sequence $P_{0},P_{1},\ldots,P_{T}$ of problems that have the aforementioned desired properties and that hence implies a lower bound for $P_{0}$. The automatic round elimination technique greatly simplifies this part, as it is able to automatically construct such a sequence. While it does not make it trivial to prove all the lower bounds that we care about (e.g., we still do not have lower bounds, as a function of $\Delta$, for $(\Delta+1)$-vertex coloring), this technique has been used to prove many interesting lower bounds for fundamental graph problems (see, e.g., [BBH+21, BO20, BBO22, BBK+21, BBK+22, BBE+20, BBK+23, BBC+25]). For example, the automatic version of round elimination made it possible to solve a long-standing open question regarding the hardness of the maximal matching problem: [^2] A matching in a graph $G$ is a selection of node-disjoint edges, and a matching in $G$ is maximal if it is not a proper subset of any other matching in $G$. indeed, [BBH+19] showed that any randomized LOCAL algorithm that solves maximal matching requires $\Omega(\min\{\Delta,\operatorname{log}_{\Delta}\operatorname{log}n\})$ rounds. Moreover, this technique heavily simplified existing proofs: it is now possible to obtain a two-line proof of the hardness of sinkless orientation.
@@
The KMW construction uses indistinguishability-type arguments to show lower bounds for many fundamental graph problems. For example, it shows that any randomized LOCAL algorithm that solves maximal matching requires $\Omega(\min\{\operatorname{log}\Delta/\operatorname{log}\operatorname{log}\Delta,\operatorname{log}_{\Delta}n\})$ rounds, even on trees. Since a maximal matching on some graph $G$ is a maximal independent set on the line graph of GG, [^3] An independent set in a graph $G$ is a selection of pairwise non-adjacent nodes, and an independent set in $G$ is maximal if it is not a proper subset of any other independent set in $G$. The line graph $G^{\prime}$ of a graph $G$ is constructed as follows: for any edge of $G$ we have a node in $G^{\prime}$ and there is an edge between two nodes in $G^{\prime}$ if and only if the edges in $G$ that they represent share a node. the KMW lower bound directly extends to maximal independent set as well, but in this specific case the lower bound does not hold on trees. Perhaps the turning point that helped to better understand the KMW lower bound happened in 2020, when [CL21] presented a breezing proof of the KMW lower bound. This paper made the KMW construction more accessible, and it helped in proving new lower bounds inspired by KMW. Indeed, shortly after, [BGK+22] came up with a novel adaptation of KMW, showing, among other things, a lower bound for maximal independent set on trees, resolving the complexity of maximal independent set on trees up to an $O(\sqrt{\operatorname{log}\operatorname{log}n})$ factor and closing an open question stated in [BE13].
@@
Very recently, [KS25] showed that any LOCAL randomized algorithm that solves maximal matching requires $\Omega(\min\{\operatorname{log}\Delta,\operatorname{log}_{\Delta}n\})$ rounds in the LOCAL model, improving the KMW lower bound (see Table 1 for a summary on the randomized lower bounds for maximal matching). This result was obtained through a novel application of the round elimination technique, called round elimination via self-reduction.
While the automatic round elimination technique of [Bra19] works for so-called locally checkable problems, this new technique works by constructing a sequence of optimization problems. In more detail, the authors of [KS25] constructed a sequence $P_{0},\ldots,P_{T}$, where each problem $P_{i}$ is the problem of approximating a maximum matching, but where the approximation factor decreases at each step. In other words, the authors show that, if we can solve in $T$ rounds a matching where the unmatched nodes are not too many, then in one round less we can produce a matching that still has a large-enough size. By repeating this reasoning iteratively, the authors get their lower bound result. This is a novel and very nice idea. However, unfortunately, this new lower bound is involved and its proof for maximal matching spans more than 25 pages. This makes the round elimination via self-reduction technique hard to digest and generalize to other graph problems.
@@
KMW $\Omega(\min\{\operatorname{log}\Delta/\operatorname{log}\operatorname{log}\Delta,\operatorname{log}_{\Delta}n\})$ $\Omega(\sqrt{\operatorname{log}n/\operatorname{log}\operatorname{log}n})$ [KMW04]
Automatic RE $\Omega(\min\{\Delta,\operatorname{log}_{\Delta}\operatorname{log}n\})$ $\Omega(\operatorname{log}\operatorname{log}n/\operatorname{log}\operatorname{log}\operatorname{log}n)$ [BBH+19]
RE via self-reduction $\Omega(\min\{\operatorname{log}\Delta,\operatorname{log}_{\Delta}n\})$ $\Omega(\sqrt{\operatorname{log}n})$ [KS25]
Table 1: The table summarizes the lower bounds for maximal matching obtained over the years. These lower bounds depend both on the maximum degree $\Delta$ and on the number $n$ of nodes of the graph. The lower bound of [BBH+19] is not comparable with the other two lower bounds: compared to the other two, this lower bound is stronger as a function of $\Delta$ and weaker as a function of $n$. The work of [KS25] strictly improves the KMW lower bound.
@@ 1.4 Our Contributions
We present a simple and clean round elimination via self-reduction technique. As a consequence, we provide a much shorter and simpler proof for obtaining the same lower bound as [KS25] for maximal matching. Our round elimination via self-reduction technique is different from the one of [KS25], and it gets rid of many new concepts that are used and needed in [KS25], resulting in a more easy-to-understand technique. We believe that this will mark a turning point for understanding the hardness of other graph problems in the LOCAL model using the round elimination via self-reduction technique. On a high level, the authors of [KS25] proved a lower bound for maximal matching by considering the sequence of problems $P_{i}$ of finding a matching where at most a fraction $p_{i}$ of nodes are unmatched, in a model where the computational entities are the edges of the graph. We show that, by considering a different problem, that we call $1$-grabbing and the more standard setting in which algorithms run on nodes, the proof becomes much simpler. This problem requires each node to select exactly one of its incident edges, and the goal is to minimize the edges grabbed by only one endpoint. In the setting considered by [KS25], the definition of the T1T-1 round algorithm requires to use the concept of $\delta$-good flowers, which complicates the definition of the algorithm and the proofs. We show that, by considering the $1$-grabbing problem, and by considering node algorithms, we can avoid this concept entirely.
@@ 2 Preliminaries
We work in the LOCAL model of distributed computing [Lin92]. In this model, we are given a distributed system of $n$ processors/nodes, that can communicate over a communication network. The network is represented as a graph $G=(V,E)$, where each node in $V$ corresponds to a processor and each edge in $E$ corresponds to a communication link between two processors. At the beginning of the computation, all nodes are identical, except for a unique identifier that is assigned to each node from the set $\{1,\ldots,n^{c}\}$ for some fixed constant $c\geq 1$, and for a local variable that stores input data. Input data consists in the input given by the considered graph problem, the degree of the node, and port numbers that allow the node to distinguish its incident communication links (note that port numbers are assigned arbitrarily by an adversary). We also assume that nodes know the number $n$ of processors in the graph. The computation proceeds in synchronous rounds and every node runs the same algorithm: in each round, a node can perform arbitrary local computation and send a message of arbitrarily large size to each of its neighbors. Computation ends when all nodes have terminated running the algorithm and output some local output. In the setting that we consider, the local output of a node is a label for each incident half-edge, which can be stored locally using port numbers. In other words, local outputs are announced on half-edges. The running time of an algorithm is the number of communication rounds until computation ends for all nodes.
@@
Randomized algorithms are allowed to fail (i.e., to not produce a correct solution for the given problem) with probability at most $1-n^{-c}$ for some constant $c\geq 1$. We say that an algorithm is no-error if it always produces a valid solution for the given problem.
@@
We will use the fact that there exist regular graphs with high girth and low independence number. The existence of such graph families has been shown in [Bol80, Bol88, BCG+22]. We use a graph family that has been already considered in other previous works (see, e.g., [BBB+24]).
Lemma 2.3 (Lemma 2.1 of [BBB+24]).
@@ 3.3 Round elimination via self-reduction
The expected badness of $\mathcal{A}_{1}$ is upper bounded by $p_{0}(1+1000\sqrt{b})$.
@@
By the definition of badness, we need to prove that $\mathbb{E}\left[E^{(1)}_{\mathsf{M}\mathsf{M}}\right]\geq b\frac{n}{2}(1-p_{0}(1+1000\sqrt{b}))$. We define $E^{(0,1)}_{\mathsf{M}\mathsf{M}}$ to be the number of edges that are labelled $\mathsf{M}\mathsf{M}$ by both $\mathcal{A}_{0}$ and $\mathcal{A}_{1}$. It holds that
@@
$\geq b\frac{n}{2}(1-p_{0})-500b^{\frac{3}{2}}np_{0}$
$=b\frac{n}{2}(1-p_{0}-1000\sqrt{b}p_{0})$
$=b\frac{n}{2}(1-(p_{0}+1000\sqrt{b}p_{0})).$
@@ 3.6 Obtaining a lower bound for trees
Clearly, maximal $b$-matching is a locally checkable problem. It is folklore that, for locally checkable problems, any lower bound for high-girth graphs implies a lower bound for trees as well (see, e.g., [Lin92, BBE+20, CKP16, BGK+22]). However, no formal proof for this generic statement has been given in previous works. Hence, for completeness, we now prove this statement.
@@ Appendix A Probabilistic inequalities
Lemma A.1 (Simplified Khintchine inequality [Haa81]).

### v2

@@ 1.1 Lower bound techniques in the LOCAL model
On a high level, lower bounds shown via the round elimination technique consist in constructing a sequence of problems that satisfy some desired properties. More in detail, let $P_{0}$ be a problem of interest for which we want to show a lower bound. The goal (and the challenge) is to produce a sequence $P_{0},P_{1},\ldots,P_{T}$ of problems such that the following two properties hold: (i) for all $0<i\leq T$ it holds that $P_{i}$ is one round easier that $P_{i-1}$; (ii) $P_{T}$ cannot be solved in 0 rounds. Such a sequence implies that $P_{0}$ cannot be solved in $T$ rounds and hence we obtain a lower bound for $P_{0}$. The first LOCAL lower bounds based on this technique date back to the late ’80s and the beginning of the ’90s, and they show that $3$-coloring cycles requires $\Omega(\operatorname{log}^{*}n)$ rounds (where $n$ is the number of nodes in the graph), both for deterministic [linial-1992-locality-in-distributed-graph-algorithms] and randomized [naor-1991-a-lower-bound-on-probabilistic-algorithms-for] algorithms. Even though at a first glance such a technique may look conceptually easy, for many years it was very challenging for researchers to use it for showing lower bounds for different graph problems in the LOCAL model. In fact, the first round elimination lower bound after the aforementioned ones was shown more than 25 years later, in a paper that presents a lower bound for the sinkless orientation problem, [^1] A solution to sinkless orientation requires orienting edges such that no node is a sink. the proof of which is very long and involved [brandt-fischer-etal-2016-a-lower-bound-for-the].
@@
Lower bounds shown via indistinguishability arguments consist in constructing a class of graphs that have some desired properties. On a very high level, this technique consists in showing that there are nodes that, within their running time, cannot distinguish between two different settings in which they would be required to give different outputs, reaching a contradiction. Perhaps one of the most influential indistinguishability-type lower bounds is the so-called KMW lower bound, shown in 2004, which provides a lower bound for many fundamental graph problems [kuhn-moscibroda-wattenhofer-2004-what-cannot-be]. The KMW construction is very nice and clever. However, the proof is quite complex. For more than 15 years, the KMW construction was one of the few (non-trivial) indistinguishability-type lower bounds, and in fact, until recently, there were no new constructions inspired by the KMW one. Arguably, the involved proof of the KMW lower bound may have been an obstacle.
@@ 1.2 Simplifying proofs and techniques marked a turning point
Regarding lower bounds proved via the round elimination technique, the turning point happened in 2019, when the automatic round elimination technique was introduced [brandt-2019-an-automatic-speedup-theorem-for]. Before this work, we had to guess what could be the sequence $P_{0},P_{1},\ldots,P_{T}$ of problems that have the aforementioned desired properties and that hence implies a lower bound for $P_{0}$. The automatic round elimination technique greatly simplifies this part, as it is able to automatically construct such a sequence. While it does not make it trivial to prove all the lower bounds that we care about (e.g., we still do not have lower bounds, as a function of $\Delta$, for $(\Delta+1)$-vertex coloring), this technique has been used to prove many interesting lower bounds for fundamental graph problems (see, e.g., [balliu-brandt-etal-2021-lower-bounds-for-maximal, brandt-olivetti-2020-truly-tight-in-delta-bounds-for, balliu-brandt-olivetti-2022-distributed-lower-bounds, balliu-brandt-etal-2021-improved-distributed-lower, balliu-brandt-etal-2022-distributed-delta-coloring, balliu-brandt-etal-2020-classification-of-distributed, balliu-brandt-etal-2023-distributed-maximal-matching, balliu-brandt-etal-2025-distributed-quantum-advantage]). For example, the automatic version of round elimination made it possible to solve a long-standing open question regarding the hardness of the maximal matching problem: [^2] A matching in a graph $G$ is a selection of node-disjoint edges, and a matching in $G$ is maximal if it is not a proper subset of any other matching in $G$. indeed, [balliu-brandt-etal-2019-lower-bounds-for-maximal] showed that any randomized LOCAL algorithm that solves maximal matching requires $\Omega(\min\{\Delta,\operatorname{log}_{\Delta}\operatorname{log}n\})$ rounds. Moreover, this technique heavily simplified existing proofs: it is now possible to obtain a two-line proof of the hardness of sinkless orientation.
@@
The KMW construction uses indistinguishability-type arguments to show lower bounds for many fundamental graph problems. For example, it shows that any randomized LOCAL algorithm that solves maximal matching requires $\Omega(\min\{\operatorname{log}\Delta/\operatorname{log}\operatorname{log}\Delta,\operatorname{log}_{\Delta}n\})$ rounds, even on trees. Since a maximal matching on some graph $G$ is a maximal independent set on the line graph of GG, [^3] An independent set in a graph $G$ is a selection of pairwise non-adjacent nodes, and an independent set in $G$ is maximal if it is not a proper subset of any other independent set in $G$. The line graph $G^{\prime}$ of a graph $G$ is constructed as follows: for any edge of $G$ we have a node in $G^{\prime}$ and there is an edge between two nodes in $G^{\prime}$ if and only if the edges in $G$ that they represent share a node. the KMW lower bound directly extends to maximal independent set as well, but in this specific case the lower bound does not hold on trees. Perhaps the turning point that helped to better understand the KMW lower bound happened in 2020, when [coupette-lenzen-2021-a-breezing-proof-of-the-kmw-bound] presented a breezing proof of the KMW lower bound. This paper made the KMW construction more accessible, and it helped in proving new lower bounds inspired by KMW. Indeed, shortly after, [balliu-ghaffari-etal-2022-node-and-edge-averaged] came up with a novel adaptation of KMW, showing, among other things, a lower bound for maximal independent set on trees, resolving the complexity of maximal independent set on trees up to an $O(\sqrt{\operatorname{log}\operatorname{log}n})$ factor and closing an open question stated in [barenboim-elkin-2013-distributed-graph-coloring].
@@
Very recently, [khoury2025] showed that any LOCAL randomized algorithm that solves maximal matching requires $\Omega(\min\{\operatorname{log}\Delta,\operatorname{log}_{\Delta}n\})$ rounds in the LOCAL model, improving the KMW lower bound (see Table 1 for a summary on the randomized lower bounds for maximal matching). This result was obtained through a novel application of the round elimination technique, called round elimination via self-reduction.
While the automatic round elimination technique of [brandt-2019-an-automatic-speedup-theorem-for] works for so-called locally checkable problems, this new technique works by constructing a sequence of optimization problems. In more detail, the authors of [khoury2025] constructed a sequence $P_{0},\ldots,P_{T}$, where each problem $P_{i}$ is the problem of approximating a maximum matching, but where the approximation factor decreases at each step. In other words, the authors show that, if we can solve in $T$ rounds a matching where the unmatched nodes are not too many, then in one round less we can produce a matching that still has a large-enough size. By repeating this reasoning iteratively, the authors get their lower bound result. This is a novel and very nice idea. However, unfortunately, this new lower bound is involved and its proof for maximal matching spans more than 25 pages. This makes the round elimination via self-reduction technique hard to digest and generalize to other graph problems.
@@
KMW $\Omega(\min\{\operatorname{log}\Delta/\operatorname{log}\operatorname{log}\Delta,\operatorname{log}_{\Delta}n\})$ $\Omega(\sqrt{\operatorname{log}n/\operatorname{log}\operatorname{log}n})$ [kuhn-moscibroda-wattenhofer-2004-what-cannot-be]
Automatic RE $\Omega(\min\{\Delta,\operatorname{log}_{\Delta}\operatorname{log}n\})$ $\Omega(\operatorname{log}\operatorname{log}n/\operatorname{log}\operatorname{log}\operatorname{log}n)$ [balliu-brandt-etal-2019-lower-bounds-for-maximal]
RE via self-reduction $\Omega(\min\{\operatorname{log}\Delta,\operatorname{log}_{\Delta}n\})$ $\Omega(\sqrt{\operatorname{log}n})$ [khoury2025]
Table 1: The table summarizes the lower bounds for maximal matching obtained over the years. These lower bounds depend both on the maximum degree $\Delta$ and on the number $n$ of nodes of the graph. The lower bound of [balliu-brandt-etal-2019-lower-bounds-for-maximal] is not comparable with the other two lower bounds: compared to the other two, this lower bound is stronger as a function of $\Delta$ and weaker as a function of $n$. The work of [khoury2025] strictly improves the KMW lower bound.
@@ 1.4 Our Contributions
We present a simple and clean round elimination via self-reduction technique. As a consequence, we provide a much shorter and simpler proof for obtaining the same lower bound as [khoury2025] for maximal matching. Our round elimination via self-reduction technique is different from the one of [khoury2025], and it gets rid of many new concepts that are used and needed in [khoury2025], resulting in a more easy-to-understand technique. We believe that this will mark a turning point for understanding the hardness of other graph problems in the LOCAL model using the round elimination via self-reduction technique. On a high level, the authors of [khoury2025] proved a lower bound for maximal matching by considering the sequence of problems $P_{i}$ of finding a matching where at most a fraction $p_{i}$ of nodes are unmatched, in a model where the computational entities are the edges of the graph. We show that, by considering a different problem, that we call $1$-grabbing and the more standard setting in which algorithms run on nodes, the proof becomes much simpler. This problem requires each node to select exactly one of its incident edges, and the goal is to minimize the edges grabbed by only one endpoint. In the setting considered by [khoury2025], the definition of the T1T-1 round algorithm requires to use the concept of $\delta$-good flowers, which complicates the definition of the algorithm and the proofs. We show that, by considering the $1$-grabbing problem, and by considering node algorithms, we can avoid this concept entirely.
@@ 2 Preliminaries
We work in the LOCAL model of distributed computing [linial-1992-locality-in-distributed-graph-algorithms]. In this model, we are given a distributed system of $n$ processors/nodes, that can communicate over a communication network. The network is represented as a graph $G=(V,E)$, where each node in $V$ corresponds to a processor and each edge in $E$ corresponds to a communication link between two processors. At the beginning of the computation, all nodes are identical, except for a unique identifier that is assigned to each node from the set $\{1,\ldots,n^{c}\}$ for some fixed constant $c\geq 1$, and for a local variable that stores input data. Input data consists in the input given by the considered graph problem, the degree of the node, and port numbers that allow the node to distinguish its incident communication links (note that port numbers are assigned arbitrarily by an adversary). We also assume that nodes know the number $n$ of processors in the graph. The computation proceeds in synchronous rounds and every node runs the same algorithm: in each round, a node can perform arbitrary local computation and send a message of arbitrarily large size to each of its neighbors. Computation ends when all nodes have terminated running the algorithm and output some local output. In the setting that we consider, the local output of a node is a label for each incident half-edge, which can be stored locally using port numbers. In other words, local outputs are announced on half-edges. The running time of an algorithm is the number of communication rounds until computation ends for all nodes.
@@
Randomized algorithms are allowed to fail (i.e., to not produce a correct solution for the given problem) with probability at most $n^{-c}$ for some constant $c\geq 1$. We say that an algorithm is no-error if it always produces a valid solution for the given problem.
@@
We will use the fact that there exist regular graphs with high girth and low independence number. The existence of such graph families has been shown in [bollobas80, bollobas88, brandt-chang-etal-2022-local-problems-on-trees-from-the]. We use a graph family that has been already considered in other previous works (see, e.g., [balliu-boudier-etal-2024-tight-lower-bounds-in-the]).
Lemma 2.3 (Lemma 2.1 of [balliu-boudier-etal-2024-tight-lower-bounds-in-the]).
@@ 3.3 Round elimination via self-reduction
The expected badness of $\mathcal{A}_{1}$ is upper bounded by $p_{0}(1+4000\sqrt{b})$.
@@
By the definition of badness, we need to prove that $\mathbb{E}\left[E^{(1)}_{\mathsf{M}\mathsf{M}}\right]\geq b\frac{n}{2}(1-p_{0}(1+4000\sqrt{b}))$. We define $E^{(0,1)}_{\mathsf{M}\mathsf{M}}$ to be the number of edges that are labelled $\mathsf{M}\mathsf{M}$ by both $\mathcal{A}_{0}$ and $\mathcal{A}_{1}$. It holds that
@@
$\geq b\frac{n}{2}(1-p_{0})-2000b^{\frac{3}{2}}np_{0}$
$=b\frac{n}{2}(1-p_{0}-4000\sqrt{b}p_{0})$
$=b\frac{n}{2}(1-(p_{0}+4000\sqrt{b}p_{0})).$
@@ 3.6 Obtaining a lower bound for trees
Clearly, maximal $b$-matching is a locally checkable problem. It is folklore that, for locally checkable problems, any lower bound for high-girth graphs implies a lower bound for trees as well (see, e.g., [linial-1992-locality-in-distributed-graph-algorithms, balliu-brandt-etal-2020-classification-of-distributed, chang-kopelowitz-pettie-2016-an-exponential-separation, balliu-ghaffari-etal-2022-node-and-edge-averaged]). However, no formal proof for this generic statement has been given in previous works. Hence, for completeness, we now prove this statement.
@@ 4 Lower bound for $(\Delta+k)$-edge coloring
Acknowledgments
We would like to thank Maxime Flin for spotting a wrong constant in a proof.
@@ Appendix A Probabilistic inequalities
Lemma A.1 (Simplified Khintchine inequality [haagerup1981]).